### v1

@@ Abstract
Tool calling has become increasingly popular for Large Language Models (LLMs). However, for large tool sets, the resulting tokens would exceed the LLM’s context window limit, making it impossible to include every tool. Hence, an external retriever is used to provide LLMs with the most relevant tools for a query. Existing retrieval models rank tools based on the similarity between a user query and a tool description (TD). This leads to suboptimal retrieval as user requests are often poorly aligned with the language of TD. To remedy the issue, we propose ToolDreamer, a framework to condition retriever models to fetch tools based on hypothetical (synthetic) TD generated using an LLM, i.e., description of tools that the LLM feels will be potentially useful for the query. The framework enables a more natural alignment between queries and tools within the language space of TD’s. We apply ToolDreamer on the ToolRet dataset and show that our method improves the performance of sparse and dense retrievers with and without training, thus showcasing its flexibility. Through our proposed framework, our aim is to offload a portion of the reasoning burden to the retriever so that the LLM may effectively handle a large collection of tools without inundating its context window.
[Uncaptioned image] ToolDreamer: Instilling LLM Reasoning Into Tool Retrievers
@@ 1 Introduction
Figure 1: Illustration of the problem statement/motivation. Retrievers conditioned to learn query-tool relationships struggle to locate tools that are not readily obvious from the question (top). LLM-generated hypothetical tools can potentially be used to locate the right tool but, it still misses the mark as it has not been conditioned to learn such tool-tool relationships (middle). Thus, if we first align the hypothetical tool with its gold tool counterpart, and incorporate the query background to train the retriever, it would lead to a far superior model capable of accurately locating the correct tool (bottom).
@@
To address the OOM issue, LLMs are paired with external retrievers Xu et al. (2025) which aim to select the right tool(s) for a query. Traditionally, tool retrievers are trained on triplets (question, gold tool, negative tool) by optimising the InfoNCE contrastive loss van den Oord et al. (2019); Henderson et al. (2017) between the three components. This means that retrieval models learn to maximise the similarity between the questions and their corresponding gold (required) tools and minimise the similarity between the questions and negative (irrelevant) tools Shi et al. (2025); Qu et al. (2024); Xu et al. (2024). However, we argue that this is a suboptimal formulation for tool retrieval and illustrate this in Fig. 1. In the given query, there is no obvious indication for weather information, which is the most relevant here, considering a BBQ is an outdoor event strongly influenced by weather conditions. Retriever models rely on semantic similarity and cues from the provided text. As such, they are unable to perform the necessary reasoning needed to select the correct tool, and consequently return an incorrect tool when using the query directly.
@@
Therefore, in this paper, we study a novel problem of retrieval optimization for retrieving tools based on hypothetical tools generated for a query. We see two challenges here, (i) data scarcity: considering the uniqueness of this problem, we do not find any existing datasets that enable training retrievers using question-hypothetical-gold tool samples, and (ii) knowledge infusion: determining the best way to integrate external knowledge as hypothetical tool descriptions into a retriever is a non-trivial task. To address these challenges, we propose a novel framework called ToolDreamer. With ToolDreamer we first generate hypothetical tools for a question through LLM-prompting. These hypothetical tools are then aligned/mapped to their gold counterpart, using bipartite graph matching, to form the training target. The aligned tools and their associated questions form the dataset used to train the retriever model, addressing the first issue. There are various ways to represent the input hypothetical tools to the retriever. ToolDreamer couples questions with their aligned tools, by integrating LLM-generated hypothetical tool metadata, to train the retriever using an augmented InfoNCE loss that utilises both question and potential tools (§3.1.3), thus tackling the second issue (Fig. 1 - bottom).
@@ 2 Related Work
Dynamic Tool Creation Recent years have seen the adoption of dynamic tool creation, i.e., using LLMs to synthesise new tools Qian et al. (2023); Yuan et al. (2024); Cai et al. (2024); Wölflein et al. (2025). This not only removes the need to write code for an extensive number of tools, but also allows for on-demand tool creation, i.e., it is not necessary to determine the entire tool-set a priori. However, the major issue with these studies is the complexity and cost involved in crafting robust code, which requires several rounds of iteration and even multi-agent processes Wölflein et al. (2025). On the other hand, our framework does not task an LLM to generate the actual implementation of the tool, but rather descriptions of tools that can potentially be used to solve that query, which is a much simpler proposition for them.
@@
Tool Retriever Training Current approaches for training retrievers Shi et al. (2025); Qu et al. (2024); Xu et al. (2024) optimise a contrastive loss objective between (query, gold tool, negative tool) triplets, which we argue is sub-optimal due to the aforementioned reasons (misalignment between tool descriptions and user queries). Our framework improves on these methods by using a better aligned anchor term, i.e., instead of training on query-tool pairs, we train on tool-tool pairs.
@@
Targeted Tool Retrieval via Auxiliary Information Generation Recently, there has been an uptick in methods for generating better search requests for tool retrieval. We identify three studies in this regard Kachuee et al. (2025); Chen et al. (2024b); Braunschweiler et al. (2025). Chen et al. (2024b) propose a training-free framework where, instead of generating tools, they generate queries from tool descriptions. These queries are then appended to the tool description to form an augmented target for test-time questions. While the tool set now contains useful additional information, it comes in the form of queries, which, as we have shown (Fig. 1), is a suboptimal target for retrieval due to their nuanced nature. Braunschweiler et al. (2025) propose a training-free ReAct Yao et al. (2023) agent that iteratively refines the query and tool set provided by the retriever. Similar to us, this framework uses an LLM to suggest potential tools for a query. However, they do not optimise their retriever using these potential tools, relying instead on several rounds of LLM reasoning, which can in turn accrue cost, particularly for API-based models. Kachuee et al. (2025) is the most similar to us in ideology. They leverage an LLM’s commonsense knowledge to generate hypothetical tool descriptions for retrieval, and also train their LLM through SFT (Supervised Fine-Tuning) to generate high-quality tools. The key difference between our methods is that they train their LLM for tool-generation, using query-tool similarity (suboptimal as explained), and ignore retriever optimisation.
Our proposed framework is inherently different from the works mentioned above. As the retriever is not optimized for matching hypothetical tool descriptions to real tools, resulting in suboptimal retrieval performance, we propose a novel framework that fine-tunes the retriever and improves the performance with better retrieval query design.
@@ 3 The Proposed ToolDreamer
Figure 2: ToolDreamer overview. Training Phase (Top): A strong LLM first generates hypothetical tools (HTs) for a question. These tools are then aligned with their gold tool (GT) counterparts. Finally, the aligned tool pairs are used to train the retrieval model. Inference Phase (Bottom): At test time, an unseen question is provided to the same LLM which generates a set of hypothetical tools for it. For each tool, the trained retriever collects a top-K list of tools from the database. A unified list is obtained through reciprocal rank fusion (RRF).
@@ 3.1 Training Phase
Our objective is to obtain a better retriever that is trained to understand user queries and relationships between LLM-generated hypothetical tools and gold tools for the associated query. However, the main challenge is obtaining the training data for the retriever, as we do not have the ground truth hypothetical-gold tool matching for each query. To address the challenge, we propose a novel two-step approach for training data generation: given a question/query, an LLM generates potentially useful HT’s. Next, we align the generated HT’s with their gold counterpart using a bipartite graph matching algorithm to form the triplets (§1). Though the alignment between HT and GT might not be perfect, it still gives a rough and sometimes accurate preference/matching between HT and GT, making it possible to train the retriever. We train our retriever to assign high similarity between the aligned tools and low similarity between HT and randomly selected tools. This equips it with ability to learn HT-GT relationships for a question, to perform a more targeted tool search.
@@ 3.1.1 Hypothetical Tool Generation
This is the first step of the training phase. Here, given a question, we prompt an LLM to generate tools it thinks could be relevant for solving the question (See Fig. 7 for Prompt Template). As we have access to the training data, we know a priori the number of tools required for the question. Thus, the LLM is instructed to generate hypothetical tools equal to the number of gold tools for the query. For the actual generation, we task the model to provide its thought process for the hypothetical tool, an appropriate name and description. The first two components are used because it has been shown Yugeswardeenoo et al. (2024); Zhu et al. (2025) that LLM performance improves when providing a reasoning breakdown before the final answer. Additionally, as mentioned, we do not generate an implementation of the tool, but only metadata (thoughts + name + description).
Although LLMs have gotten better over the years, they are still prone to making mistakes and not following instructions Tong et al. (2024); Chen et al. (2024a). We notice a similar phenomenon in our study. While our LLM (§4) is able to generate the right number of tools for the majority of questions, it falters for a very small fraction (0.3%). These questions are subsequently removed from the sampled dataset.
@@ 3.1.2 Tool Alignment
We first compute a semantic similarity (using Qwen3-8B Zhang et al. (2025b) embedding) matrix for each (hypothetical tool, gold tool) pair. From this matrix, we perform minimal cost matching using the Hungarian algorithm Kuhn (1955). The algorithm solves the problem: given a task-cost matrix, what is the optimal way of assigning tasks to minimise overall cost. The assignment is done in a way to create a best possible 1:1 mapping between elements. We leverage the same method to optimally connect tools based on their similarity and create our set of aligned (HT, GT) pairs.
A natural concern here is misalignment4.3.2) (poor HT-GT match). As we deal with a square semantic similarity matrix, there will always be a match. However, it is a best possible attempt by the underlying matching algorithm and embedding model used to score HT-GT similarities. Noise can be greatly mitigated by using a strong embedding model and a reliable alignment method. As mentioned previously (§3.1), HT-GT matching might not be perfect. However, this imperfect mapping acts as a rough, and often accurate proxy for the true alignment, making it possible to train the retriever. The ultimate goal is to train the retriever to assign high similarity between the aligned tools and low similarity between HT and random tools. We do this to make it perform a more targeted tool search, using the learned HT-GT relationships.
@@ 3.1.3 Retriever Training
Each HT is associated with three components, i.e., the thought process of the LLM behind suggesting that tool, the tool name, and its description. There are different ways to integrate this information (§1) into the retriever for selecting gold tools. We could use the HT name and description only as input to the retriever. However, the LLM-thoughts also contain important reasoning information for suggesting the necessity of the generated HT, which provides additional signal to the retriever. As such, to take advantage of the entire metadata, we format our input to the retriever as,
@@
Overall, our final loss function is given in Eq. 1. $A$ = TND / QTND is the anchor, i.e., the reference for which the model has to be optimised; $GT$ = Gold Tool; $n_{i}$ is the $i^{th}$ negative/irrelevant tool for a training sample, and sim = semantic similarity function, usually cosine similarity. $\cdot$] represents the embedding for the corresponding term.
@@ 3.2.2 Tool Retrieval
An LLM is not always able to follow instructions correctly. Consequently, for a very small fraction of questions, it does not generate equal number of thoughts, names and descriptions. As we require each HT to be associated with this metadata to form the final search vectors, to handle these instances, we ignore the HT generations and fallback on the base question as the search vector. Despite repeated prompt changes, the LLM was unable to generate the correct number of components for all questions, i.e., it was found to always falter for some questions. However, this is not a major issue as the number of such samples is not large enough to introduce significant performance degradation.
@@ 3.2.3 Retrieval Unification
For a given question, there are multiple search vectors, based on the number of generated hypothetical tools, formatted according to a particular strategy (TND/QTND). Thus, the retriever returns multiple top-K tool lists, one for each vector. As not all tools in these top-K lists are useful, we need to filter out non-important ones and unify the useful ones into a single list for the LLM. One straightforward way is to select the top-1 tool from each retrieved list. However, as these lists are generated for the same question, they might have dependencies and overlaps that such simple methods do not take into consideration. Thus, to combine each top-K retrieved tool list into a single one, we propose to use Reciprocal Rank Fusion (RRF) Cormack et al. (2009). RRF is a widely adopted technique Samuel et al. (2025); Kuo et al. (2025); Kassaie et al. (2025) to unify multiple ranked lists. It works on the principle of creating a ranked list that takes into account the relative position of items across each list. In other words, if an item in each list is consistently ranked high, it will be reflected in the final list also.
@@ 4 Experiments
In this section, we conduct experiments to evaluate the effectiveness of the proposed framework. We aim to answer three research questions: (i) the benefit/performance gains obtained by using our framework over current approaches; (ii) the relative importance of each component of our framework and (iii) whether our method is flexible, i.e., independent of the choice of underlying models.
Training
Split Web Code Customized Avg.
@@
Table 1: Performance of BM25 and Qwen3 using ToolDreamer. Best scores are in bold. Percentage improvements in teal w.r.t either zero-shot (using questions only) performance (testing only) or when training on question-tool mapping (Qwen3). Q = Trained on question-tool mapping (ToolRet (TR) baseline); TND = Thought + Name + Descriptions; QTND = Question + TND; 0s = zero-shot. The training script provided by Qu et al. (2024) for phase-2 of COLT only reports R@10/N@10.

### v2

@@ Abstract
Tool calling has become increasingly popular for Large Language Models (LLMs). However, for large tool sets, the resulting tokens would exceed the LLM’s context window limit, making it impossible to include every tool. Hence, an external retriever is used to provide LLMs with the most relevant tools for a query. Existing retrieval models rank tools based on the similarity between a user query and a tool description (TD). This leads to suboptimal retrieval as user requests are often poorly aligned with the language of TD. To remedy the issue, we propose ToolDreamer, a framework that conditions retriever models to fetch tools based on hypothetical (synthetic) TD generated using an LLM, i.e., descriptions of tools that the LLM feels will be potentially useful for the query. The framework enables a more natural alignment between queries and tools within the language space of TD’s. We apply ToolDreamer on the ToolRet dataset and show that our method improves the performance of sparse and dense retrievers with and without training, showcasing its flexibility. With our proposed framework, we aim to offload a portion of the reasoning burden to the retriever so that the LLM may effectively handle a large collection of tools without inundating its context window.
ToolDreamer: Instilling LLM Reasoning Into Tool Retrievers
@@ 1 Introduction
Figure 1: Illustration of the problem statement/motivation. Retrievers conditioned to learn query-tool relationships struggle to locate tools that are not readily obvious from the question (top). LLM-generated hypothetical tools can potentially be used to locate the right tool, but it still misses the mark as it has not been conditioned to learn such tool-to-tool relationships (middle). Thus, if we first align the hypothetical tool with its gold tool counterpart, and incorporate the query background to train the retriever, it would lead to a far superior model capable of accurately locating the correct tool (bottom).
@@
To address the OOM issue, LLMs are paired with external retrievers Xu et al. (2025) which aim to select the right tool(s) for a query. Traditionally, tool retrievers are trained on triplets (question, gold tool, negative tool) by optimising the InfoNCE contrastive loss van den Oord et al. (2019); Henderson et al. (2017) between the three components. This means that retrieval models learn to maximise the similarity between the questions and their corresponding gold (required) tools and minimise the similarity between the questions and negative (irrelevant) tools Shi et al. (2025); Qu et al. (2024); Xu et al. (2024). However, we argue that this is suboptimal for tool retrieval and illustrate this in Fig. 1 (top). In the given query, there is no obvious indication for weather information, which is the most relevant here, considering a BBQ is an outdoor event strongly influenced by weather conditions. Retriever models rely on semantic similarity and cues from the provided text. As such, they are unable to perform the necessary reasoning needed to select the correct tool, and consequently return an incorrect tool when using the query directly.
@@
Therefore, in this paper, we study a novel problem of retrieval optimization for retrieving tools based on hypothetical tools generated for a query. We see two challenges here: (i) data scarcity: considering the uniqueness of this problem, we do not find any existing datasets that enable training retrievers using question-hypothetical-gold tool samples, and (ii) knowledge infusion: determining the best way to integrate external knowledge as hypothetical tool descriptions into a retriever is a non-trivial task. To address these challenges, we propose a novel framework called ToolDreamer. With ToolDreamer, we first generate hypothetical tools for a question through LLM-prompting. These hypothetical tools are then aligned/mapped to their gold counterpart, using bipartite graph matching, to form the training target. The aligned tools and their associated questions form the dataset used to train the retriever model, addressing the first issue. There are various ways to represent the input hypothetical tools to the retriever. ToolDreamer couples questions with their aligned tools, by integrating LLM-generated hypothetical tool metadata, to train the retriever using an augmented InfoNCE loss that utilises both question and potential tools (§3.1.3), thus tackling the second issue (Fig. 1 - bottom).
@@ 2 Related Work
Dynamic Tool Creation. Recent years have seen the adoption of dynamic tool creation, i.e., using LLMs to synthesise new tools Qian et al. (2023); Yuan et al. (2024); Cai et al. (2024); Wölflein et al. (2025). This not only removes the need to write code for an extensive number of tools, but also allows for on-demand tool creation, i.e., it is not necessary to determine the entire tool-set a priori. However, the major issue with these studies is the complexity and cost involved in crafting robust code, which requires several rounds of iteration and even multi-agent processes Wölflein et al. (2025). On the other hand, our framework does not task an LLM to generate the actual implementation of the tool, but rather descriptions of tools that can potentially be used to solve that query, which is a much simpler proposition for them.
@@
Tool Retriever Training. Current approaches for training retrievers Shi et al. (2025); Qu et al. (2024); Xu et al. (2024) optimise a contrastive loss objective between (query, gold tool, negative tool) triplets, which we argue is sub-optimal due to the aforementioned reasons (misalignment between tool descriptions and user queries). Our framework improves on these methods by using a better aligned anchor term, i.e., instead of training on query-tool pairs, we train on tool-tool pairs.
@@
Tool Retrieval With Auxiliary Information. Recently, there has been an uptick in methods for generating better search requests for tool retrieval. Chen et al. (2024b) propose a training-free framework where, instead of generating tools, they generate queries from tool descriptions. These queries are then appended to the tool description to form an augmented target for test-time questions. While the tool set now contains additional useful information, it comes in the form of queries, which, as we have shown (Fig. 1), is a suboptimal target for retrieval due to their nuanced nature. Braunschweiler et al. (2025) propose a training-free ReAct Yao et al. (2023) agent that iteratively refines the query and tool set provided by the retriever. Similar to us, this framework uses an LLM to suggest potential tools for a query. However, they do not optimise their retriever using these potential tools, relying instead on several rounds of LLM reasoning, which can in turn accrue cost, particularly for API-based models. Kachuee et al. (2025) is the most similar to us in ideology. They leverage an LLM’s commonsense knowledge to generate hypothetical tool descriptions for retrieval, and also train their LLM through SFT (Supervised Fine-Tuning) to generate high-quality tools. The key difference between our methods is that they train their LLM for tool-generation, using query-tool similarity (suboptimal as explained), and ignore retriever optimisation. Our proposed framework is inherently different from these works. As the retriever is not optimized for matching hypothetical tool descriptions to real tools, resulting in suboptimal retrieval performance, we propose a novel framework that fine-tunes the retriever and improves the performance with better retrieval query design.
Instructional Retrieval. A line of work related to the above is enhancing queries with additional instructions to guide retrieval. However, these works either use manually crafted Weller et al. (2025); Sun et al. (2024) or LLM-generated instructions Shi et al. (2025) to get retrieval directives. While more informative than the original query, these instructions may not match well with the tools space or fully decompose the query with tool description language, ending up with a suboptimal alignment in tool retrieval. In contrast, our approach (i) enforces the thought process of the LLM to better match the underlying meaning of the query with tools space (Fig. 1: the weather tool is required while the query does not mention weather at all.), and (ii) appropriately decomposes a query to cover cases requiring multiple tools to answer (Fig. 4).
@@ 3 The Proposed ToolDreamer
Figure 2: ToolDreamer overview. Training Phase (Top): A strong LLM first generates hypothetical tools (HTs) for a question. These tools are then aligned with their gold tool (GT) counterparts. Finally, the aligned tool pairs are used to train the retrieval model. Inference Phase (Bottom): At test time, an unseen question is provided to the same LLM, which generates a set of hypothetical tools for it. For each tool, the trained retriever collects a top-K list of tools from the database. A unified list is obtained through reciprocal rank fusion (RRF).
@@ 3.1 Training Phase
Our objective is to obtain a better retriever that is trained to understand user queries and relationships between LLM-generated hypothetical tools and gold tools for the associated query. However, the main challenge is obtaining the training data for the retriever, as we do not have the ground truth hypothetical-to-gold tool matching for each query. To address the challenge, we propose a novel two-step approach for training data generation: given a question/query, an LLM generates potentially useful HT’s. Next, we align the generated HT’s with their gold counterpart using a bipartite graph matching algorithm to form the triplets (§1). Though the alignment between HT and GT might not be perfect, it still gives a rough and sometimes accurate preference/matching between HT and GT, making it possible to train the retriever. We train our retriever to assign high similarity between the aligned tools and low similarity between HT and randomly selected tools. This equips it with the ability to learn HT-GT relationships for a question and to perform a more targeted tool search.
@@ 3.1.1 Hypothetical Tool Generation
This is the first step of the training phase. Here, given a question, we prompt an LLM to generate tools it thinks could be relevant for solving the question (See Fig. 7 for Prompt Template). As we have access to the training data, we know a priori the number of tools required for the question. Thus, the LLM is instructed to generate hypothetical tools equal to the number of gold tools for the query. For the actual generation, we task the model to provide its thought process for the hypothetical tool, an appropriate name, and a description. The first two components are used because it has been shown Yugeswardeenoo et al. (2024); Zhu et al. (2025) that LLM performance improves when providing a reasoning breakdown before the final answer. Additionally, as mentioned, we do not generate an implementation of the tool, but only metadata (thoughts + name + description).
Although LLMs have gotten better over the years, they are still prone to making mistakes and not following instructions Tong et al. (2024); Chen et al. (2024a). We notice a similar phenomenon in our study. While our LLM (§4) can generate the right number of tools for the majority of questions, it falters for a very small fraction (0.3%). These questions are subsequently removed from the sampled dataset.
@@ 3.1.2 Tool Alignment
We first compute a semantic similarity (using Qwen3-8B Zhang et al. (2025b) embedding) matrix for each (hypothetical tool, gold tool) pair. From this matrix, we perform minimal cost matching using the Hungarian algorithm [^1] The Hungarian algorithm has a time complexity of $O(n^{3})$ where $n$ is the number of elements (rows/columns of the cost matrix) to align. This can potentially introduce significant computation overhead into the framework. However, this holds only if the cost matrix to optimize is significantly large. In ToolRet, the number of tools required for a question ranges between 1 (56% of questions) and 8 (0.1%), with the average being 2. As such, their resultant matrices do not take considerable time to process. Kuhn (1955). The algorithm solves the problem: given a task-cost matrix, what is the optimal way of assigning tasks to minimise overall cost? The assignment is done to create a best possible 1:1 mapping between elements. We leverage the same method to optimally connect tools based on their similarity and create our set of aligned (HT, GT) pairs.
A natural concern here is misalignment4.3.2) (poor HT-GT match). As we deal with a square semantic similarity matrix, there will always be a match. However, it is a best possible attempt by the underlying matching algorithm and embedding model used to score HT-GT similarities. Noise can be greatly mitigated by using a strong embedding model and a reliable alignment method. As mentioned previously (§3.1), HT-GT matching might not be perfect. However, this imperfect mapping acts as a rough and often accurate proxy for the true alignment, making it possible to train the retriever. The ultimate goal is to train the retriever to assign high similarity between the aligned tools and low similarity between HT and random tools. We do this to make it perform a more targeted tool search, using the learned HT-GT relationships.
@@ 3.1.3 Retriever Training
Each HT is associated with three components, i.e., the thought process of the LLM behind suggesting that tool, the tool name, and its description. There are different ways to integrate this information (§1) into the retriever for selecting gold tools. We could use the HT name and description only as input to the retriever. However, the LLM-thoughts also contain important reasoning information for suggesting the necessity of the generated HT, which provides an additional signal to the retriever. Thus, to take advantage of the entire metadata, we format our input to the retriever as,
@@
Overall, our final loss function is given in Eq. 1. $A$ = TND / QTND is the anchor, i.e., the reference for which the model has to be optimised; $GT$ = Gold Tool; $n_{i}$ is the $i^{th}$ negative/irrelevant tool for a training sample, and sim = semantic similarity function, usually cosine similarity. $\cdot$] represents the embedding for the corresponding term. We provide training/hardware details in App. C.
@@ 3.2.2 Tool Retrieval
An LLM is not always able to follow instructions correctly. Consequently, for a very small fraction of questions, it does not generate an equal number of thoughts, names, and descriptions. As we require each HT to be associated with this metadata to form the final search vectors, to handle these instances, we ignore the HT generations and fall back on the base question as the search vector. Despite repeated prompt changes, the LLM was unable to generate the correct number of components for all questions, i.e., it was found to always falter for some questions. However, this is not a major issue as the number of such samples is not large enough to introduce significant performance degradation.
@@ 3.2.3 Retrieval Unification
For a given question, there are multiple search vectors, based on the number of generated hypothetical tools, formatted according to a particular strategy (TND/QTND). Thus, the retriever returns multiple top-K tool lists, one for each vector. As not all tools in these top-K lists are useful, we need to filter out non-important ones and unify the useful ones into a single list for the LLM. One straightforward way is to select the top-1 tool from each retrieved list. However, as these lists are generated for the same question, they might have dependencies and overlaps that such simple methods do not take into consideration. Thus, to combine each top-K retrieved tool list into a single one, we propose using Reciprocal Rank Fusion (RRF) Cormack et al. (2009). RRF is a widely adopted technique Samuel et al. (2025); Kuo et al. (2025); Kassaie et al. (2025) to unify multiple ranked lists. It works on the principle of creating a ranked list that takes into account the relative position of items across each list. In other words, if an item in each list is consistently ranked high, it will be reflected in the final list also.
@@ 4 Experiments
In this section, we conduct experiments to evaluate the effectiveness of the proposed framework. We aim to answer four research questions (RQ): (i) the benefit/performance gains obtained by using our framework over current approaches; (ii) the relative importance of each component of our framework; (iii) whether our method is flexible, i.e., independent of the choice of underlying models and (iv) additional considerations for usage in production.
Train? Split Web Code Customized Avg.
@@
Table 1: Performance of BM25 and Qwen3 using ToolDreamer. Best scores are in bold. Percentage improvements in teal w.r.t either zero-shot (using questions only) performance (testing only) or when training on question-tool mapping (Qwen3). Q = Trained on question-tool mapping (ToolRet (TR) baseline); TND = Thought + Name + Descriptions; QTND = Question + TND; 0s = zero-shot. The training script provided by Qu et al. (2024) for phase-2 of COLT only reports R@10/N@10. For each split, we only use their associated tools when performing retrieval. Averages are computed using each split’s performance.